Sophie is taking her family to a baseball game. Adult tickets cost $12 and child tickets cost $4. She paid $64 total for the tickets. She bought 4 more child tickets than adult tickets. How many adult tickets and how many child tickets did she buy? If x represents the number of adult tickets bought , how represent the number of child tickets bought using “x”
Answer by addingup(3677)

You say: "If x represents the number of adult tickets bought... represent the number of child tickets using x"
Which is it? x = adult or x = child?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Here's your answer, A for adults and C for children. Change the one you want for an x:
C = A+4 we'll use this value for C next
12A + 4C = 64
12A + 4(A+4) = 64
12A + 4A + 16 = 64
16A = 48
A = 3 She bought 3 adult tickets
And the problem says that C = A+4 = 3 + 4 = 7 She bought 7 children tickets
===========================================
Check:
3(12) + 7(4) = 36 + 28 = 64 Correct